OpenStudy (anonymous):

Explain, in complete sentences, how you would use the elimination method to solve the following system of equations and provide the solution to the system. 5x - 3y = 4 3x - 5y = -4

OpenStudy (anonymous):

What would you do first?

OpenStudy (anonymous):

I Dont Know!

OpenStudy (anonymous):

start by solving for x in the 1st equation

OpenStudy (anonymous):

how do you get rid of the -3y?

OpenStudy (anonymous):

@meanzie

OpenStudy (anonymous):

:) You want to eliminate one of the variables by adding the two equations together like this 5x - 3y = 4 3x - 5y = -4 --------------- 8x - 8y = 0 This particular addition doesn't give you any useful information, because we haven't eliminated a variable. How do you think you could alter one or both of these equations it so that adding them together leaves you with either x or y by itself?

OpenStudy (anonymous):

Well It Doesent Matter Now I Ran Outta Time On My Test :(

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

better to visit before the test